OpenStudy (anonymous):

solve the equation for x. Write the exact solution and the approximate solution to the nearest hundredth, when approppriate. 9^(x) = 6

OpenStudy (anonymous):

alright so you know how to use the logarithmic function to move ^x ?

OpenStudy (anonymous):

in other words if you multiply both sides by Ln Ln (9^x)= Ln 6

OpenStudy (anonymous):

(x)(ln (9))=ln(6)

OpenStudy (anonymous):

divide by ln 9 to get

OpenStudy (anonymous):

\[\frac{\ln(6)}{\ln(9)}\]

OpenStudy (anonymous):

= x

OpenStudy (anonymous):

thank youuuuu

OpenStudy (anonymous):

no problem =]